The half-life of a certain radioactive isotope is 32 hours. What fraction of the sample would remain after 16 huors? a. 0.50 b. 0.25 c. 0.62 d. 0.71
1
Expert's answer
2017-02-21T08:30:43-0500
The formula for radioactive decay is

N=N0 * (1/2)^(t/tau)

where N0 is initial quantity of isotope and tau is half-life.
so we find
N/N0 = (1/2)^(t/tau) = (1/2)^(1/2) = 0.71

Answer is d. 0.71